# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ IV-A3 Speech sources
The semantics in speeches, which encompass not only the vocabulary and grammar but also the intonation, timbre, and speed of speech, are significantly complex and subjective. The same sentence with different intonations can convey different meanings. There are two evaluation methods, respectively evaluating the meaning similarity of the texts contained in the speeches or the speeches themselves. For the former, we can use all the aforementioned metrics for text sources. For the latter, we focus on the “meaning similarity” of speeches, which can be interpreted as the similarity degree of the inferred meanings derived from the speeches at the transmitter and receiver. The metrics related to the second method are introduced below. Objective metrics. Assume that the original and recovered speeches are respectively and . Similar to the SNR, the signal-to-distortion ratio (SDR) [81] was initially proposed to calculate the ratio between the signal quality and distortion. As an extension, the SDR can be used to measure the similarity between two speeches[86].
@@
where represents the energy of . A larger SDR indicates a higher similarity between speeches. Another metric called perceptual evaluation of speech quality (PESQ)[82] calculates the speech quality score by comparing the differences between speeches, which was established as Recommendation ITU-T P.862 in 2001. The specific process of calculating the PESQ score for two speeches mainly includes two stages: the alignment routine and the perceptual model. The former consists of the computation of the overall system gain, the intelligent reflecting surface (IRS) filtering, the time alignment (including the envelope-based alignment, the fine time alignment, the utterance splitting, and the perceptual realignment), and so on. The latter includes the IRS-receive filtering, the short-term fast fourier transform (FFT), the calculation of the pitch power densities, the realignment of bad intervals, the computation of the PESQ score, and so on. The range of the PESQ score is -0.5 to 4.5, and the larger value reveals the higher similarity between speeches. The calculation process of the PESQ is quite complex, and due to limited space, we have provided only a brief description here. For more details, see, e.g., [82] and the references therein. Furthermore, the Mel cepstral distortion (MCD)[83], which measures the spectral distortion between synthesized and target speeches, can be regarded as a metric to evaluate the similarity between speeches. The smaller the MCD is, the closer the spectra between speeches are. The aforementioned metrics do not sufficiently represent the semantic information of speeches, making it difficult to accurately evaluate the similarity between speeches at the semantic level. Learning-based metrics. Based on the FID and KID, the unconditional Fréchet DeepSpeech distance (FDSD) and unconditional kernel DeepSpeech distance (KDSD) are proposed in [84] to measure the meaning similarity between speeches. By utilizing the pre-trained DeepSpeech2 model from the NVIDIA OpenSeq2Seq library, the KDSD and FDSD are able to capture the semantic information in the high-level speech feature space. The computational differences between the FDSD and FID are only reflected in the source and network types. According to (23), the FDSD is defined as
@@
where , are the mean vector and covariance matrix in the feature space of the original speech distribution, and , are the ones of the recovered speech distribution. indicates the trace of a square matrix. Moreover, the KDSD is given by
@@
where , are respectively the number of samples in the original and recovered speeches. , are the vectors obtained from the DeepSpeech2 model for the original speech, and , are the ones for the recovered speech, and is the polynomial kernel defined as (25). Additionally, [84] also proposes the conditional Fréchet DeepSpeech distance (cFDSD) and conditional kernel DeepSpeech distance (cKDSD) to evaluate the meaning similarity between speeches under the specific linguistic feature distribution condition. Compared to objective metrics, learning-based metrics utilize NNs to assist in capturing the semantic information, which aid in better representing the meaning similarity of speeches.
@@ IV-B Methods: Based on DNNs
Shannon’s separation theorem[12] proves that under the condition of infinite block lengths, the separate source-channel coding (SSCC) scheme can achieve the same performance as the JSCC scheme. Due to its modularity and ease of implementation, the SSCC scheme is widely applied in traditional communication systems, with source codes and channel codes being proposed in parallel on separate tracks. However, affected by the typical characteristics of large-scale scenarios, highly dynamic channels, and limited device capabilities in the SAGSIN, the communication systems have extremely stringent requirements on latency, while the computing resources of IoT devices are limited. Moreover, in the semantic communication systems, redundant information is removed through the semantic extraction (a kind of perception) before encoding. All these facts lead to the prevalence of finite block lengths (even short ones) in the SAGSIN. Thus, the SSCC scheme cannot achieve the Shannon limit, and the JSCC one regains attention. The mainstream approach in constructing semantic communication systems currently involves utilizing DNNs as source (semantic) and channel codecs, and training them jointly, called deep learning based JSCC (DeepJSCC), which is first applied in [89] to implement semantic communications. Taking the transmitter as an example here, DNNs sequentially perform operations such as semantic extraction and encoding on the sampled data. The output of the network can be directly transmitted into the channel. Therefore, it is necessary to review typical DNNs which can be used in JSCC. In semantic communication systems, commonly used DNN-related models primarily include four categories: convolutional neural Networks (CNNs), recurrent neural networks (RNNs), Transformer and generative adversarial networks (GANs). These networks become the base of the system in Section IV-C.
@@ IV-B1 CNN-related models
CNN is primarily used for space relevant sources, particularly image ones. The development of CNN can be traced back to 1962 when Hubel and Wiesel introduced the concept of Receptive fields [90]. In 1980, [91] proposed a neural network including convolutional and pooling layers. The first formal CNN model, called LeNet-5[92], was proposed in 1998. It laid the foundation for the basic structure of CNNs: convolutional, pooling, and fully connected layers. However, due to limited computing resources and the vanishing gradient caused by sigmoidal activation functions, LeNet-5 did not receive the deserving attention. AlexNet [93] marks the triumphant return of CNN, which consists of eight layers, including five convolutional ones and three fully connected ones. Compared to LeNet-5, AlexNet has the following four main improvements: First, a deeper network architecture is used to learn more intricate semantic features. Second, the sigmoidal activation function is replaced by the rectified linear unit (ReLU) one, alleviating the vanishing gradient. Third, researchers utilize two data augmentation methods and dropout technique to reduce overfitting. Fourth, overlapping pooling further enhances the richness of feature maps and extracts more detailed semantic information. Later, ZFNet[94] (a fine-tuned version of AlexNet), VGGNet[95] (a “very deep” version of AlexNet), and GoogLeNet [96] (also known as Inception-V1) are successively proposed. The aforementioned CNNs are primarily used for image classification tasks, and thus the network output is the classification result. However, for image semantic segmentation tasks, the network should output an image of the same size as the input one and classify each pixel of the image into semantic categories (e.g., person, car, and tree). Based on this purpose, [97] proposes the fully convolutional network (FCN), which consists entirely of convolutional layers.
The development history of CNNs implies a belief: increasing the network depth does not decrease the model performance. A straightforward example is that adding a layer with the “identity mapping” does not change the model performance, but makes the network deeper. However, “Is learning better networks as easy as stacking more layers?[98]” Researchers discovered the “degradation” phenomenon through experiments, revealing a negative answer to the question. In order to make the network fully learn the data semantic features, activation functions are set to be nonlinear, to the extent that deep layers cannot simply perform the linear transformation (e.g., “identity mapping”). One solution is to build a “shortcut connection” between certain layers to strike better balance between nonlinear and linear transformations, forming a residual block. This is the ResNet[98]. Recently, some new variants of CNNs have been also proposed consecutively, including DenseNet[99] (employing dense connection structures), PSPNet[100] (utilizing pyramid pooling), and PatchGAN[101] (commonly used as a discriminator in GANs).
@@ IV-B2 RNN-related models
RNN is mainly used for time relevant sources, particularly text and speech ones. In 1982, a Hopfield network was introduced by [102] as a precursor to RNNs. Later, a Jordan network was proposed in [103]. In [104], a Jordan network was improved, resulting in a Elman network, which is also known as the standard RNN. It consists of an input layer, a hidden layer, and an output layer. The neurons in the hidden layer are recurrently connected over time, allowing the output from the previous time step to be fed as input to the current one, thus enabling the network to remember and utilize historical information. This is why RNNs can process sequential data and learn semantic features from the context. When the input sequence is too long, the time steps of the standard RNN become large, which may lead to the vanishing/exploding gradients during training. The solution is to utilize techniques like gradient clipping or more stable RNN variants such as long short-term memory (LSTM) networks. LSTM[105] effectively addresses the issue of long-term dependencies in standard RNNs by introducing memory units and gate mechanisms (an input gate, a forget gate and an output gate). It can efficiently propagate and express information in long sequences without neglecting/forgetting useful information from earlier time steps. In fact, only a portion of the input sequence is crucial (i.e., having more semantic information), and it needs to be remembered for a long time, while the remaining one can be appropriately forgotten. Therefore, the concept of “long short-term memory” was proposed in [105]. Because standard RNNs only consider past information of the input sequence, [106] introduced two independent RNN modules (forward and backward processing) simultaneously to capture both past and future context information, resulting in the bidirectional RNN (BiRNN). Afterwards, [107] combined the essence of LSTM and BiRNN to create bidirectional LSTM (BiLSTM) networks. Furthermore, [108] proposed deep BiLSTM networks to achieve better extraction and representation of semantic features. Gated recurrent unit (GRU) networks[109], a variant of LSTM, controls the flow of information by using an update gate and a reset gate. It is simpler than the LSTM model and can achieve comparable performance to LSTM with higher training efficiency.
@@ IV-B3 Transformer-related models
The Transformer is powerful due to the following aspects: Firstly, the multi-head self-attention mechanism divides the model into multiple heads to form subspaces and allows the model to automatically learn the relationships between different positions in the input sequence, capturing contextual information more effectively. Secondly, the self-attention mechanism allows for “parallel” computation of the encoder and decoder, effectively harnessing the parallel computing capabilities of modern hardware accelerators such as graphics processing units (GPUs). Thirdly, the Transformer effectively models long-term dependencies in sequences, alleviating the issue of gradient vanishing/exploding in RNNs when processing long sequences. Fourthly, the Transformer structure is simple, highly modular, and easy to extend and modify. In addition, there are several variants of the Transformer proposed recently, such as the vision Transformer (ViT) [111] (for image classification tasks), the detection Transformer (DETR)[112] (for object detection tasks), the shifted window (swin) Transformer[113] (reducing computational complexities by the window self-attention mechanism), and so on.
@@ IV-B4 GAN-related models
In practical applications, the GAN can be applied to various types of sources. The classic GAN was first proposed in [114], which consists of a generator and a discriminator. The purpose of the generator is to generate instances that appear natural and realistic, closely resembling the original data. And the role of the discriminator is to determine whether a given instance is real or fake. Clearly, they have an ”adversarial” relationship. The generator and discriminator are trained alternately, progressing together in the “game” and ideally reaching a dynamic equilibrium (i.e., Nash equilibrium). At this point, the generated instances from the generator are indistinguishable from real ones. There are two key points to note: Firstly, the ultimate goal is to obtain a well-trained generator, while the discriminator is an additional benefit. Secondly, there are no specific constraints on DNNs used for the generator and discriminator, and thus they can be CNNs, RNNs, or even Transformers. The initiall GAN is an unsupervised learning model, and its training process does not require labeled data. Conditional GANs (CGANs)[115] bring GANs back into the realm of supervised learning, alleviating the problem of unstable training in GANs. Deep convolutional GANs (DCGANs)[116] are the first successful implementation of GANs that use CNNs instead of fully connected layers. This significantly reduces the number of network parameters while greatly improving the quality of generated data. Wasserstein GANs (WGANs)[117] improve the stability of model training by introducing the Wasserstein distance as a loss function. Self-attention GANs (SAGANs)[118] introduce an attention mechanism that enables the network to better understand the semantic information. The introduction of BigGANs[119] reveals that GAN network training can also benefit from large-scale data and computing resources. BigGANs achieve unprecedented levels of image generation quality. Recently, some new variants of GANs have been proposed, such as StyleGANs[120] (controlling the style and features of generated images), HiFi-GAN [121] (focusing on generating high-fidelity audio signals), and so on.
@@
There are also other neural networks used in JSCC schemes. Graph convolutional networks (GCNs)[122], which are commonly utilized to extract semantic features from non-Euclidean structured images (e.g., social networks and knowledge graphs), are used in [123] to construct joint source-channel codecs. [124, 125, 87, 88, 126] employs reinforcement learning networks (RLNs) to construct semantic communication systems.

### v3

@@ IV-A3 Speech sources
The semantics in speeches, which encompass not only the vocabulary and grammar but also the intonation, timbre, and speed of speech, are significantly complex and subjective. The same sentence with different intonations can convey different meanings. There are two evaluation methods, respectively evaluating the meaning similarity of the texts contained in the speeches or the speeches themselves. For the former, we can use all the aforementioned metrics for text sources. For the latter, we focus on the “meaning similarity” of speeches, which can be interpreted as the similarity degree of the inferred meanings derived from the speeches at the transmitter and receiver. The metrics related to the second method are introduced below.
Objective metrics. Assume that the original and recovered speeches are respectively and . Similar to the SNR, the signal-to-distortion ratio (SDR) [95] was initially proposed to calculate the ratio between the signal quality and distortion. As an extension, the SDR can be used to measure the similarity between two speeches[100].
where represents the energy of . A larger SDR indicates a higher similarity between speeches.
@@
Another metric called perceptual evaluation of speech quality (PESQ)[96] calculates the speech quality score by comparing the differences between speeches, which was established as Recommendation ITU-T P.862 in 2001. The specific process of calculating the PESQ score for two speeches mainly includes two stages: the alignment routine and the perceptual model. The former consists of the computation of the overall system gain, the intelligent reflecting surface (IRS) filtering, the time alignment (including the envelope-based alignment, the fine time alignment, the utterance splitting, and the perceptual realignment), and so on. The latter includes the IRS-receive filtering, the short-term fast Fourier transform (FFT), the calculation of the pitch power densities, the realignment of bad intervals, the computation of the PESQ score, and so on. The range of the PESQ score is -0.5 to 4.5, and the larger value reveals the higher similarity between speeches. The calculation process of the PESQ is quite complex, and due to limited space, we have provided only a brief description here. For more details, see, e.g., [96] and the references therein.
Furthermore, the Mel cepstral distortion (MCD)[97], which measures the spectral distortion between synthesized and target speeches, can be regarded as a metric to evaluate the similarity between speeches. The smaller the MCD is, the closer the spectra between speeches are.
The aforementioned metrics do not sufficiently represent the semantic information of speeches, making it difficult to accurately evaluate the similarity between speeches at the semantic level.
Learning-based metrics. Based on the FID and KID, the unconditional Fréchet DeepSpeech distance (FDSD) and unconditional kernel DeepSpeech distance (KDSD) are proposed in [98] to measure the meaning similarity between speeches. By utilizing the pre-trained DeepSpeech2 model from the NVIDIA OpenSeq2Seq library, the KDSD and FDSD are able to capture the semantic information in the high-level speech feature space. The computational differences between the FDSD and FID are only reflected in the source and network types. According to (23), the FDSD is defined as
@@
where , are the mean vector and covariance matrix in the feature space of the original speech distribution, , are the ones of the recovered speech distribution, and indicates the trace of a square matrix. Moreover, the KDSD is given by
@@
where , are respectively the number of samples in the original and recovered speeches, , are the vectors obtained from the DeepSpeech2 model for the original speech, , are the ones for the recovered speech, and is the polynomial kernel defined as (25). Additionally, [98] also proposes the conditional Fréchet DeepSpeech distance (cFDSD) and conditional kernel DeepSpeech distance (cKDSD) to evaluate the meaning similarity between speeches under the specific linguistic feature distribution condition.
Compared to objective metrics, learning-based metrics utilize NNs to assist in capturing the semantic information, which can better represent the meaning similarity of speeches.
@@ IV-B Methods: Based on DNNs
Shannon’s separation theorem[9] proves that under the condition of infinite block lengths, the separate source-channel coding (SSCC) scheme can achieve the same performance as the JSCC scheme. Due to its modularity and ease of implementation, the SSCC scheme is widely applied in traditional communication systems, with source codes and channel codes being proposed in parallel on separate tracks.
However, affected by the typical characteristics of large-scale scenarios, highly dynamic channels, and limited device capabilities in the SAGSIN, the communication systems have extremely stringent requirements on latency, while the computing resources of IoT devices are limited. Moreover, in the semantic communication systems, redundant information is removed through semantic extraction before encoding. All these facts lead to the prevalence of finite block lengths (even short ones) in the SAGSIN. Thus, the SSCC scheme cannot achieve the Shannon limit, and the JSCC one regains attention.
The mainstream approach in constructing semantic communication systems currently involves utilizing DNNs as source (semantic) and channel codecs, and training them jointly, called deep learning based JSCC (DeepJSCC), which is firstly applied in [103] to implement semantic communications. Taking the transmitter as an example here, DNNs sequentially perform operations such as semantic extraction and encoding on the source data. The output of the network can be directly transmitted over the channel. Therefore, it is necessary to review typical DNNs which can be used in JSCC.
In semantic communication systems, commonly used DNN-related models primarily include four categories: convolutional neural Networks (CNNs), recurrent neural networks (RNNs), Transformer and GANs. These networks, which are listed in Table VI, become the base of the system in Section IV-C.
Types | Neural Networks | Explanation and Application Scenarios |
CNN-Related | LeNet-5 | First CNN model, laying the foundation for CNN structure | [104] | |||
AlexNet | Deeper CNN, having improvements compared to LeNet-5 | [105] | ||||
FCN | Consisting of only convolutional layers for semantic segmentation | [106] | ||||
ResNet | A residual block via “shortcut connection” | [107] | ||||
Others |
RNN-Related | Hopfield/Jordan network | Initial exploration on standard RNNs | [114, 115] | |||
Elman network | Standard RNN that can utilize historical information | [116] | ||||
LSTM network | Addressing long-term dependency issue by memory units | [117] | ||||
BiRNN/BiLSTM network | Two independent networks to capture more context information | [118, 119, 120] | ||||
GRU network | Simpler than LSTM model | [121] | ||||
Transformer | Classic Transformer | Initially for NLP tasks, extracting long-term dependency | [122] | |||
Others |
[123, 124, 125] | ||||
GAN | Classic GAN | Generator-discriminator unsupervised adversarial training | [126] | |||
CGAN | Introducing supervised learning to avoid unstable training | [127] | ||||
DCGAN | Lighter GAN model implementation by CNN | [128] | ||||
WGAN | Introducing Wasserstein distance as loss function | [129] | ||||
SAGAN | Attention-based GAN | [130] | ||||
BigGAN | Large scale training for better image quality | [131] | ||||
Others |
[132, 133] |
@@ IV-B1 CNN-related models
CNN is primarily used for space relevant sources, particularly image ones. The development of CNN can be traced back to 1962 when Hubel and Wiesel introduced the concept of Receptive fields [134]. In 1980, [135] proposed a neural network including convolutional and pooling layers. The first formal CNN model, called LeNet-5[104], was proposed in 1998. It laid the foundation for the basic structure of CNNs: convolutional, pooling, and fully connected layers. However, due to limited computing resources and the vanishing gradient caused by sigmoidal activation functions, LeNet-5 did not receive the deserving attention.
AlexNet [105] marks the triumphant return of CNN, which consists of eight layers, including five convolutional ones and three fully connected ones. Compared to LeNet-5, AlexNet has the following four main improvements: First, a deeper network architecture is used to learn more intricate semantic features. Second, the sigmoidal activation function is replaced by the rectified linear unit (ReLU) one, alleviating the vanishing gradient. Third, researchers utilize two data augmentation methods and dropout technique to reduce overfitting. Fourth, overlapping pooling further enhances the richness of feature maps and extracts more detailed semantic information. Later, ZFNet[108] (a fine-tuned version of AlexNet), VGGNet[109] (a “very deep” version of AlexNet), and GoogLeNet [110] (also known as Inception-V1) are successively proposed.
The aforementioned CNNs are primarily used for image classification tasks, and thus the network output is the classification result. However, for image semantic segmentation tasks, the network should output an image of the same size as the input one and classify each pixel of the image into semantic categories (e.g., person, car, and tree). Based on this purpose, [106] proposes the fully convolutional network (FCN), which consists entirely of convolutional layers.
The development history of CNNs implies a belief: increasing the network depth does not decrease the model performance. A straightforward example is that adding a layer with the “identity mapping” does not change the model performance, but makes the network deeper. However, “Is learning better networks as easy as stacking more layers?[107]” Researchers discovered the “degradation” phenomenon through experiments, revealing a negative answer to the question. In order to make the network fully learn the data semantic features, activation functions are set to be nonlinear, to the extent that deep layers cannot simply perform the linear transformation (e.g., “identity mapping”). One solution is to build a “shortcut connection” between certain layers to strike better balance between nonlinear and linear transformations, forming a residual block. This is the ResNet[107].
@@ IV-B2 RNN-related models
RNN is mainly used for time relevant sources, particularly text and speech ones. In 1982, a Hopfield network was introduced by [114] as a precursor to RNNs. Later, a Jordan network was proposed in [115]. In [116], a Jordan network was improved, resulting in a Elman network, which is also known as the standard RNN. It consists of an input layer, a hidden layer, and an output layer. The neurons in the hidden layer are recurrently connected over time, allowing the output from the previous time step to be fed as input to the current one, thus enabling the network to remember and utilize historical information. This is why RNNs can process sequential data and learn semantic features from the context.
When the input sequence is too long, the time steps of the standard RNN become large, which may lead to the vanishing/exploding gradients during training. The solution is to utilize techniques like gradient clipping or more stable RNN variants such as long short-term memory (LSTM) networks. LSTM[117] effectively addresses the issue of long-term dependencies in standard RNNs by introducing memory units and gate mechanisms (an input gate, a forget gate and an output gate). It can efficiently propagate and express information in long sequences without neglecting/forgetting useful information from earlier time steps. In fact, only a portion of the input sequence is crucial (i.e., having more semantic information), and it needs to be remembered for a long time, while the remaining one can be appropriately forgotten. Therefore, the concept of “long short-term memory” was proposed in [117].
Because standard RNNs only consider past information of the input sequence, [118] introduced two independent RNN modules (forward and backward processing) simultaneously to capture both past and future context information, resulting in the bidirectional RNN (BiRNN). Afterwards, [119] combined the essence of LSTM and BiRNN to create bidirectional LSTM (BiLSTM) networks. Furthermore, [120] proposed deep BiLSTM networks to achieve better extraction and representation of semantic features. Gated recurrent unit (GRU) networks[121], a variant of LSTM, controls the flow of information by using an update gate and a reset gate. It is simpler than the LSTM model and can achieve comparable performance to LSTM with higher training efficiency.
@@ IV-B3 Transformer-related models
The Transformer is powerful due to the following aspects: Firstly, the multi-head self-attention mechanism divides the model into multiple heads to form subspaces and allows the model to automatically learn the relationships between different positions in the input sequence, capturing contextual information more effectively. Secondly, the self-attention mechanism allows for “parallel” computation of the encoder and decoder, effectively harnessing the parallel computing capabilities of modern hardware accelerators such as graphics processing units (GPUs). Thirdly, the Transformer effectively models long-term dependencies in sequences, alleviating the issue of gradient vanishing/exploding in RNNs when processing long sequences. Fourthly, the Transformer structure is simple, highly modular, and easy to extend and modify.
In addition, there are several variants of the Transformer proposed recently, such as the vision Transformer (ViT) [123] (for image classification tasks), the detection Transformer (DETR)[124] (for object detection tasks), the shifted window (swin) Transformer[125] (reducing computational complexities by the window self-attention mechanism), and so on.
@@ IV-B4 GAN-related models
In practical applications, the GAN can be applied to various types of sources. The classic GAN was first proposed in [126], which consists of a generator and a discriminator. The purpose of the generator is to generate instances that appear natural and realistic, closely resembling the original data. And the role of the discriminator is to determine whether a given instance is real or fake. Clearly, they have an “adversarial” relationship. The generator and discriminator are trained alternately, progressing together in the “game” and ideally reaching a dynamic equilibrium (i.e., Nash equilibrium). At this point, the generated instances from the generator are indistinguishable from real ones. There are two key points to note: Firstly, the ultimate goal is to obtain a well-trained generator, while the discriminator is an additional benefit. Secondly, there are no specific constraints on DNNs used for the generator and discriminator, and thus they can be CNNs, RNNs, or even Transformers.
The initiall GAN is an unsupervised learning model, and its training process does not require labeled data. Conditional GANs (CGANs)[127] bring GANs back into the realm of supervised learning, alleviating the problem of unstable training in GANs. Deep convolutional GANs (DCGANs)[128] are the first successful implementation of GANs that use CNNs instead of fully connected layers. This significantly reduces the number of network parameters while greatly improving the quality of generated data. Wasserstein GANs (WGANs)[129] improve the stability of model training by introducing the Wasserstein distance as a loss function. Self-attention GANs (SAGANs)[130] introduce an attention mechanism that enables the network to better understand the semantic information. The introduction of BigGANs[131] reveals that GAN network training can also benefit from large-scale data and computing resources. BigGANs achieve unprecedented levels of image generation quality. Recently, some new variants of GANs have been proposed, such as StyleGANs[132] (controlling the style and features of generated images), HiFi-GAN [133] (focusing on generating high-fidelity audio signals), and so on.
@@
There are also other neural networks used in JSCC schemes. Graph convolutional networks (GCNs)[136], which are commonly utilized to extract semantic features from non-Euclidean structured images (e.g., social networks and knowledge graphs), are used in [137] to construct joint source-channel codecs. [138, 139, 101, 102, 140] employ reinforcement learning networks (RLNs) to construct semantic communication systems.